The table below shows information about the closed economy.

Real GDP (RM million) Aggregate Expenditure (RM million) 

850 880

900 920

950 960

1000 1000

1050 1040

a) If the full employment level of income is RM1, 350 million,

i) state whether the above economy is facing an inflationary gap or deflationary gap. Give your reason.

ii) what is the effect of the gap that you named in [a(i)] to the above economy.

iii) calculate the required aggregate expenditure changes needed to eliminate the inflationary gap or deflationary gap.

b) If the full employment level of income is RM850 million,

i) state whether the above economy is facing an inflationary gap or deflationary gap. Give your reason.

ii) what is the effect of the gap that you named in [b(i)] to the above economy.

iii) calculate the required aggregate expenditure changes needed to eliminate the inflationary gap or deflationary gap.


1
Expert's answer
2021-03-26T08:05:34-0400

a.

(i) Inflationary gap. This is because real GDP is less than the potential GDP.

(ii) Inflationary gap leads to general increase of prices. When it exists at full employment, money income to people increases but output doesn't.

(iii) 1,350-850=500

1,350-900=450

1,350-950=400

1,350-1000=350

1,350-1050=300

b.

(i) Deflationary gap. This is because real GDP is greater than potential GDP.

(ii) The effect of Deflationary gap are: low rates or negative economic growth, negative effect on government's budget and lower economic growth. Government will receive lower tax revenue subsequently lower government expenditure.

(iii) 850-850=0

850-900=50

850-950=100

850-1000=150

850-1050=200